A computer can generate random numbers which are either 0 or 2. On a particular occasion, it generates a set of numbers which consists of 23 zeros and 17 twos. Find the mean and variance of this set of 40 numbers. 



1
Expert's answer
2021-02-24T06:01:07-0500
"mean=\bar{x}=\dfrac{\sum_ix_i}{n}"

"=\dfrac{23(0)+17(2)}{40}=0.85"

"Var(X)=\dfrac{\sum_i(x_i-\bar{x})^2}{n}"

"Var(X)=\dfrac{23(0-0.85)^2+17(2-0.85)^2}{40}"

"=0.9775"